# thanks: equal contributionthanks: equal contribution

### v1

@@ II Biopolymers & cytoskeletal filaments
Perhaps the most important example of microtubule self-organisation is the mitotic spindle, which partitions the set of chromosomes from a cell into two copies for the daughter cells Wittmann et al. (2001); Gadde and Heald (2004). The polarity of microtubules is essential in this process, as each half of the spindle contains uniformly oriented microtubules, with their minus-ends at the pole (towards the centrosome) and their plus-ends extending away. The initial growth of microtubules around centrosomes is based on their dynamic instability, with random phases of growth and shrinkage probing the cytoplasm for kinetochores (the so-called ‘search-and-capture’ model Kirschner and Mitchison (1986)). That many classes of kinesins play a role in the spindle formation is evidence that the mitotic spindle is a nonequilibrium self-organised structure Gadde and Heald (2004).
@@
$\bullet$ Actin filaments are composed of (globular) G-actin monomers which readily assemble into long strings called F-actin (“F” for filamentous) of about 8 nm in width Boal (2012). Actin filaments are about 300 times less stiff than microtubules, with $\beta_{\mathrm{f}}=7.3\times 10^{-26}$ N m2 Gittes et al. (1993), a persistence length $\lambda_{\mathrm{p}}\approx 18~{}\mu$m Gittes et al. (1993), and with a Young’s modulus $E\simeq(1-3)$ GPa, which is comparable with Young’s modulus for bone ($9$ GPa) Mofrad (2009). Interestingly, actin is less resistant to bending than microtubules only because it is thinner, and not because it is more compliant than tubulin Gittes et al. (1993). It is also worth comparing the Young’s moduli of actin and tubulin. The measurements of $\beta_{\mathrm{f}}$ for a microtubule implies a Young’s modulus $E\sim 1.2$ GPa, while for actin $E\sim 2.6$ GPa Gittes et al. (1993). This is comparable with $E\sim 2-8$ GPa of silk (Bombyx mori) Lepore et al. (2016); Greco et al. (2022); Wainwright (1982), or with $E\sim 2.5$ GPa of keratin (from birds’ feathers) Bonser and Purslow (1995). These are all proteins employed by the cells to generate tensile strength Gittes et al. (1993), so it is not so surprising after all that filaments with similar Young’s modulus evolved for the cytoskeleton.
@@
Both actin filaments and microtubules are polar filaments, that is, they have a positive and a negative end due to an asymmetry at the molecular level. Hence, they can function as train tracks along which molecular motors move. Dynein and kinesin are the motor proteins associated to microtubules; by means of ATP hydrolysis they can transport vesicles, mRNA, and other proteins. Myosins are the motor proteins associated to actin. They are a superfamily Hartman and Spudich (2012): at least 35 classes of motor proteins associated to actin filaments are known Odronitz and Kollmar (2007) and used for vesicle transport; they are responsible for muscle contraction in myocytes Sweeney and Houdusse (2010), and are macromolecular filaments. Some members of this superfamily, the non-muscle myosin II (NM II) are even involved in cell adhesion and migration Vicente-Manzanares et al. (2009).
@@
Figure 2: (a) Endothelial cells from cow lung with two components of the cytoskeleton: microtubules (green) and actin filaments (red) Tina Carvalho (2011). (b) Confocal microscope image of a mouse 3T3 fibroblast cell. The nucleus has been stained blue, and two components of the cytoskeleton, actin microfilaments and intermediate filaments, are stained red and green respectively David Becker (2011).
@@
Actin is present in high concentrations in the actomyosin cortex of animal cells, a layer attached to the inner surface of the cell membrane, and its function is to generate tension and regulate force and cell shape Étienne et al. (2015); Saha et al. (2016); Kumar (2021). Actin is the main constituent of lamellipodia, two-dimensional structures responsible for movement of e.g. epithelial cells Cardamone et al. (2011), and is involved in cell adhesion Schwarz and Safran (2013); Garcia et al. (2015). The versatility of actin is but an example of the fact that the molecular organisation of biofilaments is extremely flexible; as a further example, amyloid fibrils can change bending rigidity by four orders of magnitude due to varying backbone interactions in the cross-$\beta$ structure Knowles et al. (2007).
@@
Actin plays a major role during embryonic development, as the cells have to undergo a specific sequence of physical and biological changes Heisenberg and Bellaïche (2013). Physically, cell cortical forces, cell adhesion and motility together produce self-organisation both at the tissue level and within single cells, e.g., reorientation of the polarity plane Aigouy et al. (2010). Another physical factor involved in tissue morphogenesis is actin-myoosin flows (or cortical flow), that is, a cell-scale internal flow of actin components of the cortex Bray and White (1988). Various aspects are at play here: the character of the actin-myosin flow dynamics (pulsatile versus continuous) and direction (centripetal or anisotropic) influence force generation; additionally, to transmit forces to other cells, this flow needs to be coupled to adhesion complexes at the cell membrane Heisenberg and Bellaïche (2013). Although much has been understood about force generation in cells and tissue, there are still many open questions about how the exquisitely complex self-organised process is controlled by the cell.
@@
$\bullet$ Intermediate filaments (IFs) are the most flexible of the three classes of cytoskeletal filaments, but they are also a very diverse class of filaments Köster et al. (2015). Estimates of their persistence length range from $\lambda_{p}=415\pm 45$ nm for human wild-type vimentin Schopferer et al. (2009) to $\lambda_{p}=2.1\pm 0.1~{}\mu$m Nöding and Köster (2012). There are six classes of IFs Parry and Steinert (1992), which include, e.g., keratin, desmin, vimentin, and nuclear lamin. Structurally, IFs exhibit a hierarchical structure Chernyatina et al. (2015). The alpha-helical domains form a coiled-coil structural motif, composed of two parallel in-register chains. Two chains aggregate in parallel to generate an IF molecule, and then two such molecules assemble in an antiparallel fashion to form a four-chain structural unit Parry and Steinert (1992). IF proteins are highly charged, thus representing versatile polyampholytes with multiple functions due to their strong self-assembling ability Herrmann and Aebi (2016). IF is built as a complex network of ionic interactions formed by the acidic rod and the basic head domain, and, unlike microtubules and actin filaments, IFs are non-polar, and lack enzymatic activity Herrmann and Aebi (2016). IFs form networks of filaments with important mechanical properties: upon being mechanically stressed, the network becomes stiffer Janmey et al. (1991); Wen and Janmey (2011); Herrmann and Aebi (2016).
@@
Cytoskeletal networks and viscoelastic properties. The viscoelastic properties of cytoskeletal filaments are closely linked to their biologic function Mofrad (2009), because their extreme aspect ratio and rod-like structure dominate the rheological behaviour of the cell, and changes in their structure may cause gel-sol transitions observed when cells are activated or begin to move Janmey et al. (1991). Measurements of the storage modulus $G^{\prime}$ for F-actin and fibrin suggest that the molecular basis of the viscoelasticity of these two types of networks is similar and may be approximated by the steric hindrance of diffusion of long, interpenetrated filamentous polymers Janmey et al. (1991). While the viscoelastic measurements of these biofilaments can be explained with isotropic models, there is ample evidence that microtubules and F-actin undergo a transition from isotropic to liquid crystalline solutions under the influence of shear stresses Kerst et al. (1990), other macromolecules Suzuki et al. (1989), or even spontaneously Hitt et al. (1990); Viamontes et al. (2006); Hess and Ross (2017), because of the strong thermodynamic drive for long rod-like filaments to form liquid-crystalline arrays Onsager (1949); Flory (1956).
@@
Cytoskeletal filaments play a fundamental role in the mechanical properties and the biological function of the cell. This functional importance is due to not just the physical properties of isolated filaments, but, perhaps more crucially, to their structural organisation. Most cytoskeletal filaments form networks, and by controlling their architecture and connectivity cells achieve remarkable material properties Wagner et al. (2006); Gardel et al. (2008).
@@
Actin filaments are assembled by actin-binding proteins into branched, cross-linked networks; the list of such proteins is vast but includes, e.g., $\alpha$-actinin, $\beta$-spectrin, filamin, and fimbrin. These complex, dynamically-reassembling, dendritic networks generate force at the cell periphery that creates membrane protrusions (lamellipodia), that in turn produce the cell’s directed motility, essential in processes such as embryonic development, wound healing, and the immune response Pollard and Borisy (2003).
@@
The networks assembled with semi-flexible cytoskeletal filaments have unusual viscoelastic properties Pegoraro et al. (2017); Storm et al. (2005); Wagner et al. (2006); Gardel et al. (2008); Chaubet et al. (2020); Lorenz and Köster (2022). Upon increasing stress, networks of actin filaments show stress stiffening behaviour up to a critical stress of $\sigma_{c}\sim 270$ Pa, after which the networks enters a stress-softening regime that is reversible, that is, if the stress is reduced no hysteresis is observed Chaudhuri et al. (2007). A microscopic mechanism that explains such behaviour is based on the initial resistance to stretching due to entropic effects in the biofilaments, leading to the stress-stiffening regime. Above $\sigma_{c}$, some filaments under compressive stresses will buckle, which decreases the elasticity of the network and leads to the stress-softening regime Chaudhuri et al. (2007). This is at variance with the behaviour of flexible polymers networks, which exhibit stress softening only due to failure and is not reversible Mofrad (2009). Furthermore, cytoskeletal filaments show negative normal stresses, that is, when sheared they tend to pull inward Janmey et al. (2007). On the other hand, flexible-polymer networks exhibit positive normal stress. These nonlinear viscoelastic features of cytoskeletal filaments contribute to generate the structural integrity and strength of the cell.
@@
Biofilaments exist also outside of the cell. In animal cells, collagen is the most important filament found in the extracellular matrix Sun (2021). There are many types of collagen molecules, but e.g. in humans more than $90\%$ are type I collagen, found in skin and tendons. Type I collagen forms fibers in a hierarchical structures. The amino acids chains form together a triple helix, called tropocollagen, with a diameter of $\approx 1.5$ nm. Multiple tropocollagens bundle together into protofibrils (diameter $\approx 6-25$ nm). Multiple protofibrils further bundle together and form collagen fibril (up to $1\mu$m in length, $\approx 100$ nm in diameter). The fibrils can then form thicker fibers, which self-organize into a disordered network, the collagen matrix Sun (2021). The collagen matrix exhibits nonlinear elasticity, with an elastic modulus depending on concentration, pH, and ionic strength, and ranging from few Pa up to hundreds of MPa Sheu et al. (2001); Dutov et al. (2016).
@@
In plant cells, the most abundant and important filament is cellulose. This is a linear polysaccharide chain of glucose monomers. The numerous hydrogen bonds within and between chains gives them considerable stiffness and generates robust fibrils, which are used in the cell walls of plant cells. Depending on its microstructure, the Young’s modulus of cellulose ranges from $\approx 25$ GPa Eichhorn and Young (2001) (for microcrystalline cellulose) to $114$ GPa (for bacterial cellulose) Hsieh et al. (2008) or the classic measurement $138$ GPa Nishino et al. (1995).
@@
A review of cell biopolymers cannot be complete without mentioning nucleic acids: DNA and RNA. The physics of the DNA helix, or generally of nucleic acids, is surprisingly complex Strzelecka et al. (1988), and is inspiring a number of applications in nanotechnology Ramsay (1998); Seeman (2010); Saccà and Niemeyer (2012); Seeman and Sleiman (2017); Praetorius et al. (2017); Engelen and Dietz (2021); but, unfortunately, we do not have the space to address it here. There are, however, numerous excellent reviews on DNA’s and RNA’s structure and biochemical interactions Kornyshev et al. (2007); Wong and Pollack (2010); Teif and Bohinc (2011); Travers and Muskhelishvili (2015); Michieletto and Gilbert (2019) and also on application of topology to DNA Chevizovich et al. (2020); Michieletto (2024)

### v2

@@ II Biopolymers & cytoskeletal filaments
Perhaps the most important example of microtubule self-organisation is the mitotic spindle, which partitions the set of chromosomes from a cell into two copies for the daughter cells [121, 122]. The polarity of microtubules is essential in this process, as each half of the spindle contains uniformly oriented microtubules, with their minus-ends at the pole (towards the centrosome) and their plus-ends extending away. The initial growth of microtubules around centrosomes is based on their dynamic instability, with random phases of growth and shrinkage probing the cytoplasm for kinetochores (the so-called ‘search-and-capture’ model [123]). That many classes of kinesins play a role in the spindle formation is evidence that the mitotic spindle is a nonequilibrium self-organised structure [122].
Microtubules and actin filaments strongly interact. And this interaction is responsible for regulating cell shape and polarity [124]. Microtubules are anchored to actin filaments, which maintain the position of the microtubule aster [125]. Additionally, very recent work reports evidence that microtubules may serve as sensors and signalling platforms that guide the organization of the cell around the centrosome [126].
@@
$\bullet$ Actin filaments are composed of (globular) G-actin monomers which readily assemble into long strings called F-actin (“F” for filamentous) of about 8 nm in width [70]. Actin filaments are about 300 times less stiff than microtubules, with $\beta_{\mathrm{f}}=7.3\times 10^{-26}$ N m2 [72], a persistence length $\lambda_{\mathrm{p}}\approx 18~{}\mu$m [72], and with a Young’s modulus $E\simeq(1-3)$ GPa, which is comparable with Young’s modulus for bone ($9$ GPa) [71]. Interestingly, actin is less resistant to bending than microtubules only because it is thinner, and not because it is more compliant than tubulin [72]. It is also worth comparing the Young’s moduli of actin and tubulin. The measurements of $\beta_{\mathrm{f}}$ for a microtubule implies a Young’s modulus $E\sim 1.2$ GPa, while for actin $E\sim 2.6$ GPa [72]. This is comparable with $E\sim 2-8$ GPa of silk (Bombyx mori) [127, 128, 129], or with $E\sim 2.5$ GPa of keratin (from birds’ feathers) [130]. Interestingly, the values of these Young’s moduli are of the same order of magnitude. The persistence length $\lambda_{\mathrm{p}}$ of actin is of the order of $10-20~{}\mu$m and its contour length is about the same [131, 72, 132]. Thus, the ratio $\lambda_{\mathrm{p}}/L\approx 1$, making actin a semiflexible polymer. Because of these physical properties, actin exhibits many diverse functions in the cell: from cell movement to tension generation [35].
@@
Both actin filaments and microtubules are polar filaments, that is, they have a positive and a negative end due to an asymmetry at the molecular level. Hence, they can function as train tracks along which molecular motors move. Dynein and kinesin are the motor proteins associated to microtubules; by means of ATP hydrolysis they can transport vesicles, mRNA, and other proteins. Myosins are the motor proteins associated to actin. They are a superfamily [133]: at least 35 classes of motor proteins associated to actin filaments are known [134] and used for vesicle transport; they are responsible for muscle contraction in myocytes [135], and are macromolecular filaments. Some members of this superfamily, the non-muscle myosin II (NM II) are even involved in cell adhesion and migration [136].
@@
Figure 2: Cytoskeletal filaments. (a) Endothelial cells from cow lung with two components of the cytoskeleton: microtubules (green) and actin filaments (red) [137]. (b) Confocal microscope image of a mouse 3T3 fibroblast cell. The nucleus has been stained blue, and two components of the cytoskeleton, actin microfilaments and intermediate filaments, are stained red and green respectively [138].
@@
Actin is present in high concentrations in the actomyosin cortex of animal cells, a layer attached to the inner surface of the cell membrane, and its function is to generate tension and regulate force and cell shape [139, 140, 141]. Actin is the main constituent of lamellipodia, two-dimensional structures responsible for movement of e.g. epithelial cells [142], and is involved in cell adhesion [143, 144]. The versatility of actin is but an example of the fact that the molecular organisation of biofilaments is extremely flexible; as a further example, amyloid fibrils can change bending rigidity by four orders of magnitude due to varying backbone interactions in the cross-$\beta$ structure [145].
@@
Actin plays a major role during embryonic development, as the cells have to undergo a specific sequence of physical and biological changes [146]. Physically, cell cortical forces, cell adhesion and motility together produce self-organisation both at the tissue level and within single cells, e.g., reorientation of the polarity plane [147]. Another physical factor involved in tissue morphogenesis is actin-myosin flows (or cortical flow), that is, a cell-scale internal flow of actin components of the cortex [148]. Various aspects are at play here: the character of the actin-myosin flow dynamics (pulsatile versus continuous) and direction (centripetal or anisotropic) influence force generation; additionally, to transmit forces to other cells, this flow needs to be coupled to adhesion complexes at the cell membrane [146]. Although much has been understood about force generation in cells and tissue, there are still many open questions about how the exquisitely complex self-organised process is controlled by the cell.
$\bullet$ Intermediate filaments (IFs) are the most flexible of the three classes of cytoskeletal filaments, but they are also a very diverse class of filaments [149]. Estimates of their persistence length range from $\lambda_{\mathrm{p}}=415\pm 45$ nm [74] to $\lambda_{\mathrm{p}}=2.1\pm 0.1~{}\mu$m [150] (both studies for human vimentin). The typical contour lengths for IFs are much larger than $\lambda_{\mathrm{p}}$. Thus, $\lambda_{\mathrm{p}}/L\ll 1$, which makes IFs flexible polymers. They are the least studied of the cytoskeletal filaments. Their main function is to provide mechanical strength to cells, but they appear to be involved in cell shape and motility and the positioning of organelles [151].
@@
There are six classes of IFs [152], which include, e.g., keratin, desmin, vimentin, and nuclear lamin. Structurally, IFs exhibit a hierarchical structure [153]. The alpha-helical domains form a coiled-coil structural motif, composed of two parallel in-register chains. Two chains aggregate in parallel to generate an IF molecule, and then two such molecules assemble in an antiparallel fashion to form a four-chain structural unit [152]. IF proteins are highly charged, thus representing versatile polyampholytes with multiple functions due to their strong self-assembling ability [154]. IFs are built as complex networks of ionic interactions formed by the acidic rod and the basic head domain, and, unlike microtubules and actin filaments, IFs are non-polar, and lack enzymatic activity [154]. IFs form networks of filaments with important mechanical properties: upon being mechanically stressed, the network becomes stiffer [69, 155, 154].
@@
Cytoskeletal networks and viscoelastic properties. The viscoelastic properties of cytoskeletal filaments are closely linked to their biologic function [71, 156], because their extreme aspect ratio and rod-like structure dominate the rheological behaviour of the cell, and changes in their structure may cause gel-sol transitions observed when cells are activated or begin to move [69]. Measurements of the storage modulus $G^{\prime}$ for F-actin and fibrin suggest that the molecular basis of the viscoelasticity of these two types of networks is similar and may be approximated by the steric hindrance of diffusion of long, interpenetrated filamentous polymers [69]. While the viscoelastic measurements of these biofilaments can be explained with isotropic models, there is ample evidence that microtubules and F-actin undergo a transition from isotropic to liquid crystalline solutions under the influence of shear stresses [157], other macromolecules [158], or even spontaneously [159, 160, 161], because of the strong thermodynamic drive for long rod-like filaments to form liquid-crystalline arrays [162, 163].
@@
Cytoskeletal filaments play a fundamental role in the mechanical properties and the biological function of the cell. This functional importance is due to not just the physical properties of isolated filaments, but, perhaps more crucially, to their structural organisation. Most cytoskeletal filaments form networks, and by controlling their architecture and connectivity cells achieve remarkable material properties [164, 165].
@@
Actin filaments are assembled by actin-binding proteins into branched, cross-linked networks; the list of such proteins is vast but includes, e.g., $\alpha$-actinin, $\beta$-spectrin, filamin, and fimbrin. These complex, dynamically-reassembling, dendritic networks generate force at the cell periphery that creates membrane protrusions (lamellipodia), that in turn produce the cell’s directed motility, essential in processes such as embryonic development, wound healing, and the immune response [166].
Generally, the elasticity of semi-flexible polymers can have two separate origins: 1. entropic elasticity, which is due to the reduction in the number of available microstates when a filament is stretched vis-a-vis in an equilibrium state subject to thermal fluctuations; 2. enthalpic elasticity, which is caused by the change in the equilibrium distance of the molecular bonds in a filament, such as when a filament is deformed (enthalpic elasticity does not depend on thermal fluctuations). The specific details of the network will dictate the relative weight of these two types of elasticity.
The worm-like chain model can be used to derive an important result in the rheology of cytoskeletal filaments [167]. Considering a filament under strong tension $f$ aligned with the $z$-axis, there will only be small fluctuations of the tangent vector $\bm{t}\equiv(t_{x},t_{y},t_{z})$ around $\hat{\bm{z}}$. We can then write $t_{z}=1-\bm{t}_{\perp}^{2}/2+\mathcal{O}(\bm{t}_{\perp}^{4})$. The bending energy is then
[…]
$\frac{E_{b}}{k_{\mathrm{B}}T}=\frac{1}{2}\int\left[\lambda_{\mathrm{p}}q^{2}+f\right]\tilde{\bm{t}}_{\perp}^{2}\frac{dq}{2\pi}-fL\,.$ (9)
From Eq. (9) and from equipartition, one can derive the extension $z_{\mathrm{e}}$ [167]. The relation between $z_{\mathrm{e}}$ and the tension $f$ can be approximated very well by the simple formula [168]
$\frac{f\lambda_{\mathrm{p}}}{k_{\mathrm{B}}T}=\frac{z_{\mathrm{e}}}{L}+\frac{1}{4(1-z_{\mathrm{e}}/L)^{2}}-\frac{1}{4}\,.$ (10)
We note that differentiating $f$ in Eq. (10) with respect to $z_{\mathrm{e}}$ gives $f^{\prime}(z_{\mathrm{e}})\sim f^{3/2}$. While this result was derived for single semiflexible polymers [167, 169], in a biopolymer network the differential modulus $K^{\prime}\equiv\partial\sigma/\partial\gamma_{\mathrm{s}}$ measures the nonlinear change of the elastic modulus with the applied shear stress $\sigma$, where $\gamma_{\mathrm{s}}$ is the strain. In fact, the prediction $K^{\prime}\sim\sigma^{3/2}$ is observed in networks of cross-linked and bundled actin filaments [23] and in IF networks [170]. This phenomenon is termed entropic stiffening [171], because the elastic modulus grows with shear stress.
@@
The networks assembled with semi-flexible cytoskeletal filaments have unusual viscoelastic properties [51, 172, 164, 165, 173, 174]. Upon increasing stress, networks of actin filaments show stress-stiffening behaviour up to a critical stress of $\sigma_{c}\sim 270$ Pa, after which the networks enters a stress-softening regime that is reversible, that is, if the stress is reduced no hysteresis is observed [175]. A microscopic mechanism that explains such behaviour is based on the initial resistance to stretching due to entropic elasticity in the biofilament network, leading to the stress-stiffening regime. Above $\sigma_{c}$, some filaments under compressive stresses will buckle, which decreases the elasticity of the network and leads to the stress-softening regime [175]. This is at variance with the behaviour of flexible polymers networks, which exhibit stress softening only due to failure and is not reversible [71]. Furthermore, cytoskeletal filaments show negative normal stresses, that is, when sheared they tend to pull inward [176]. A negative normal stress is unusual (when a typical material is sheared, it expands in the direction normal to the shear, i.e., it experiences a positive normal stress), but it can be rationalised with a simple model of nonlinear elasticity [172, 176]. Such filament exhibits strain-stiffening behaviour: it resists extension more strongly as the extension grows. In a random, isotropic network of such filaments under shear, some filaments will be compressed while others will be stretched. The stretched filaments will exert a stronger tensile force (negative stress) than the compressed filaments (positive stress); hence, the net stress in the normal direction will be negative [177, 172, 176]. This is to be contrasted with flexible-polymer networks, which exhibit positive normal stress. The nonlinear viscoelastic features of cytoskeletal filaments contribute to generate the structural integrity and strength of the cell.
@@
Biofilaments exist also outside of the cell. In animal cells, collagen is the most important filament found in the extracellular matrix [178]. There are many types of collagen molecules, but e.g. in humans more than $90\%$ are type I collagen, found in skin and tendons. Type I collagen forms fibres in a hierarchical structures. The amino acids chains form together a triple helix, called tropocollagen, with a diameter of $\approx 1.5$ nm. Multiple tropocollagens bundle together into protofibrils (diameter $\approx 6-25$ nm). Multiple protofibrils further bundle together and form collagen fibril (up to $1\mu$m in length, $\approx 100$ nm in diameter). The fibrils can then form thicker fibres, which self-organize into a disordered network, the collagen matrix [178]. The collagen matrix exhibits nonlinear elasticity, with an elastic modulus depending on concentration, pH, and ionic strength, and ranging from few Pa up to hundreds of MPa [179, 180].
@@
In plant cells, the most abundant and important filament is cellulose. This is a linear polysaccharide chain of glucose monomers. The numerous hydrogen bonds within and between chains gives them considerable stiffness and generates robust fibrils, which are used in the cell walls of plant cells. Depending on its microstructure, the Young’s modulus of cellulose ranges from $\approx 25$ GPa [181] (for microcrystalline cellulose) to $114$ GPa (for bacterial cellulose) [182] or the classic measurement $138$ GPa [80].
@@
A review of cell biopolymers cannot be complete without mentioning nucleic acids: DNA and RNA. The physics of the DNA helix, or generally of nucleic acids, is surprisingly complex [183], and is inspiring a number of applications in nanotechnology [184, 185, 186, 187, 188, 189]; but, unfortunately, we do not have the space to address it here. There are, however, numerous excellent reviews on DNA’s and RNA’s structure and biochemical interactions [190, 191, 192, 193, 194] and also on application of topology to DNA [195, 196]
The cytoskeleton is a formidable subject to study theoretically or computationally, because of the sheer number of individual units involved and their complex physical and biochemical interactions. However, considerable progress has been made recently [197, 198, 199], and software packages are now available that simulate cytoskeletal structures of various shapes and their dynamics [200, 201, 202, 203, 204]. Often cytoskeletal filaments are modelled as rigid rods, however, a versatile model marries ideas from polymer physics and active matter to simulate active filaments [205, 206, 68, 207, 208, 209, 210, 208, 211]. Such model polymers are commonly represented as a semi-flexible chain of $N$ colloids connected by springs. In an overdamped regime, the equations of motion for the monomer positions $\bm{r}_{\alpha}$ read
$\gamma\frac{d\bm{r}_{\alpha}}{dt}=-\nabla_{\alpha}U+\bm{F}_{\mathrm{a},\alpha}+\bm{\xi}_{\alpha}(t)\,,$ (11)
where $\gamma$ is the monomers’ drag coefficient, $U=U_{\text{bond}}+U_{\text{angle}}+U_{\alpha,\beta}$ the configurational potential energy, with contributions from the bonds between neighbouring monomers
$U_{\text{bond}}=\frac{k}{2}\sum_{\alpha=1}^{N-1}\left(|\bm{r}_{\alpha}-\bm{r}_{\alpha+1}|-r_{0}\right)^{2}\,\text{,}$ (12)
with spring constant $k$ and equilibrium bond length $r_{0}$, from a bending energy
$U_{\text{angle}}=\frac{\kappa}{4}\sum_{\alpha=1}^{N-2}\left(2\bm{r}_{\alpha+1}-\bm{r}_{\alpha}-\bm{r}_{\alpha+2}\right)^{2}\,\text{,}$ (13)
with bending rigidity $\kappa$ and from pairwise steric interactions between monomers $U_{\alpha,\beta}$ commonly modelled as a Weeks–Chandler–Andersen potential [212]. In Eq. (11), $\bm{\xi}(t)$ is a white noise term describing thermal fluctuations with $\langle\xi_{\alpha,i}(t)\xi_{\beta,j}(t^{\prime})\rangle=2\gamma k_{\text{B}}T\delta(t-t^{\prime})\delta_{i,j}\delta_{\alpha,\beta}$. The active force term $\bm{F}_{\mathrm{a}_{\alpha}}$ drives the system out of equilibrium. A common choice is a tangential driving force $\bm{F}_{\mathrm{a},\alpha}=f(\bm{r}_{\alpha+1}-\bm{r}_{\alpha})$, however, depending on the filament’s driving mechanism other choice of $\bm{F}_{\mathrm{a},\alpha}$ are possible. This model has been successfully applied to larger systems, such as flagella [213], cyanobacteria [214], and worms [215].
A different approach to the theoretical modelling of cytoskeletal filaments is the active nematic framework. In vitro assays of microtubules driven by kinesin motors can self-organise into arrays of asters and vortices [216]. Addition of a depleting agent, leads to the formation of bundles in a quasi-2D layer of microtubules and molecular motors, which produce chaotic flows [217]. This system has been the experimental archetype for the theory of active nematics. The theory considers a coarse-grained order parameter that describes the degree of nematic order, that is, the degree of alignment between filaments with head-tail symmetry. A local director $\bm{n}$, with nematic symmetry $\bm{n}\leftrightarrow-\bm{n}$ reflects the average local filament alignment direction and a second-rank, traceless tensor is the order parameter $\mathbf{Q}=\frac{d}{d-1}S(\bm{n}\otimes\bm{n}-\mathbf{I}/d)$, where $d$ is the dimensionality of space, $q$ the magnitude of nematic order, and $\mathbf{I}$ the identity tensor. The dynamics of the active gel is governed by the nematodynamics of $\mathbf{Q}$ and the incompressible Navier–Stokes equation for the flow field $\bm{u}$
$\frac{\partial\mathbf{Q}}{\partial t}+\bm{u}\cdot\nabla\mathbf{Q}-\mathbf{S}=\Gamma\mathbf{H}\,,$ (14)
$\rho\left(\frac{\partial\bm{u}}{\partial t}+\bm{u}\cdot\nabla\bm{u}\right)=\nabla\cdot\mathbf{\Pi}\,,\quad\nabla\cdot\bm{u}=0\,,$ (15)
where $\mathbf{S}$ is the corotation term, $\mathbf{H}$ the molecular field, and $\mathbf{\Pi}$ is the stress tensor, which includes the hydrostatic pressure, the viscous and elastic contributions, and, crucially, the active stress tensor $\mathbf{\Pi}^{\mathrm{act}}=-\zeta\mathbf{Q}$ (see, e.g. [218] for the full definitions of $\mathbf{S}$, $\mathbf{H}$, and $\mathbf{\Pi}$). The coupling of the flow field with the derivative of the nematic tensor $\mathbf{Q}$ generates the rich phenomenology of active nematics and their topological defects [219, 220, 221, 222, 223, 224, 225, 226].
Box III | Drag anisotropy Drag anisotropy refers to the directional dependence of drag forces acting on an object as it moves through a medium. Due to their elongated form, filaments experience less hydrodynamic drag when pulled along their length than when pulled perpendicular to it. Drag anisotropy can be quantified as the ratio $\xi=\frac{\zeta_{\perp}}{\zeta_{\parallel}},$ (16) where $\zeta_{\perp}$ and $\zeta_{\parallel}$ are the friction coefficients for motion perpendicular and parallel to the object’s length. Typically, for filaments $\zeta_{\perp}\approx 2\zeta_{\parallel}$. The total drag force density (drag per unit length) opposing a filament’s motion at low Reynolds number (i.e. viscosity-dominated regime) is given by $\textbf{F}_{\text{d}}=-(\zeta_{\parallel}\textbf{v}_{\parallel}+\zeta_{\perp}\textbf{v}_{\perp})$, where $\textbf{v}_{\parallel}$ and $\textbf{v}_{\perp}$ are the projections of the local velocity into directions parallel and perpendicular to the segment [11, 227].